OpenStudy (anonymous):

The ratio of the base edges of two similar pyramids is 3:4. The volume of the larger pyramid is 320 in3. What is the volume of the smaller pyramid? 189 in3 180 in3 240 in3 135 in3

OpenStudy (anonymous):

A pyramids volume is \[\frac{ lwh }{ 3 }\] Where l, w are the side lengths and h is the height

OpenStudy (anonymous):

Now in this question I assume both pyramids have same height but the edges differ in a ration to 3:4 So both l, w changes. This the actual area ratio is 3^2 : 4^2 or 9:16 320/16=20 9*20=180